Kutztown University Department of Mathematics

Colloquium

11:00 a.m.

April 2, 2009

Lytle Hall 228

The Topology of Links and Knots

(or, how I learned to enjoy math)

Dr. Valentina Joukhovitski

Assistant Professor of Mathematics Department of Mathematics University of Michigan

Abstract

We will discuss knots and some of their invariants. These will tell us whether we can untangle a knot, or not. We will start with elementary polynomial invariants and, if we are lucky, eventually will graduate to scarier sounding words.

11:00 a.m. talk begins

Kutztown Univ. of Pennsylvania Dept. of Mathematics Dr. Paul Ache, Chair